# GlitchMiner: Mining Glitch Tokens in Large Language Models via Gradient-based Discrete Optimization

### v4

@@ 4.2 Main Results
These experimental findings underscore GlitchMiner’s ability not only to accurately detect glitch tokens but also to generalize across a diverse set of model architectures. This adaptability and precision position GlitchMiner as a powerful tool for enhancing the robustness and security of LLMs.
@@ 4.4 Token Entropy Analysis
These findings reinforce the effectiveness of GlitchMiner’s entropy-based optimization in efficiently detecting glitch tokens by focusing on areas of high prediction uncertainty within the model.
@@ 5 Conclusion
In this paper, we introduced GlitchMiner, a novel framework for detecting glitch tokens in LLMs through gradient-based discrete optimization. Our method combines entropy-based loss functions with a local search strategy to navigate token space efficiently, improving both detection accuracy and computational efficiency. Experimental results across diverse LLMs highlight GlitchMiner’s robustness and versatility compared to existing methods. Our findings confirm GlitchMiner’s effectiveness in identifying glitch tokens across architectures, demonstrating its potential for enhancing the security and reliability of LLM-based applications. Future work could explore extending this approach to other high-dimensional discrete optimization tasks.
Broader Impact
The potential impact of this research extends beyond mere academic interest. By improving the detection of glitch tokens, GlitchMiner contributes to enhancing the overall reliability and safety of LLMs in real-world applications. This is particularly crucial in high-stakes domains such as healthcare, finance, and autonomous systems, where the consequences of model errors can be severe. Furthermore, our approach opens new avenues for understanding and mitigating vulnerabilities in AI systems, potentially influencing the development of more robust and trustworthy AI technologies in the future.
@@ Acknowledgments
We would like to express our sincere gratitude to Sander Land for his valuable insights and feedback on this work, particularly regarding the evaluation metrics. His suggestions were instrumental in refining our approach and improving the clarity of our results. We also wish to thank Guanlin Li for his valuable suggestions on our manuscript. His insights were crucial for improving the content and structure of the paper.

### v5

@@ 4.1 Experimental Setup
Baselines. We compare our proposed glitch token detection method with two state-of-the-art approaches: GlitchHunter (GlitchHunter) and Magikarp (Magikarp). These methods serve as the primary benchmarks for evaluating our approach.
Although GlitchProber (GlitchProber) is another relevant method, it follows a fundamentally different approach by pre-collecting a subset of glitch tokens to train a classifier, introducing a supervised learning component. In contrast, GlitchMiner, along with GlitchHunter and Magikarp, uses heuristic-based methods to detect glitch tokens without relying on labeled data or additional classifier training. This methodological difference makes a direct comparison less meaningful, so we focus our evaluation on methods that align more closely with our unsupervised approach.
Parameter Settings. In our implementation of GlitchMiner, we use KK=32 and BB=8 as the default parameters. These values were chosen based on empirical testing to balance computational efficiency and detection effectiveness. Specifically, KK=32 defines the size of the local neighborhood considered in each iteration, while BB=8 determines the batch size for entropy computation. These settings have shown to provide a good trade-off between exploration of the token space and exploitation of local information across various model architectures.
4.2 Main Results
The performance of GlitchMiner against the baselines is detailed in Table 2. The results provide strong evidence for the effectiveness of our entropy-guided search, showing that GlitchMiner consistently discovers more glitch tokens than both GlitchHunter and Magikarp under fixed query budgets.
On average, GlitchMiner outperforms the strongest baseline, Magikarp, by 10.7% on the Detected@2000 metric. This performance advantage is not uniform but reveals an important trend: while Magikarp’s simple norm-based heuristic can be effective for an initial, low-budget scan (e.g., on Llama-3.1-8B), GlitchMiner’s more sophisticated local search strategy consistently proves more fruitful as the search budget expands. For instance, on Llama-2-7B-chat-hf, GlitchMiner finds nearly three times as many glitches as Magikarp (532 vs. 186), demonstrating its superior ability to uncover less obvious candidates that simple heuristics miss.
These findings validate that framing glitch detection as a behavior-driven optimization problem is a more robust and generalizable strategy than relying on static, model-specific patterns.
@@ 4.4 Token Entropy Analysis
These findings validate the effectiveness of GlitchMiner’s entropy-based optimization in efficiently detecting glitch tokens by focusing on areas of high prediction uncertainty within the model.
@@ 5 Conclusion
In this paper, we introduced GlitchMiner, a novel framework that detects glitch tokens in LLM by reframing the problem as a behavior-driven optimization task. Departing from prior work that relies on static patterns, our approach identifies anomalous tokens by searching for inputs that maximize the model’s predictive uncertainty, measured by output entropy. We operationalize this principle with an efficient gradient-guided local search strategy that can accurately pinpoint high-entropy candidates in the vast token space. Extensive experiments on 10 diverse LLMs demonstrated that GlitchMiner significantly outperforms state-of-the-art baselines in both detection accuracy and efficiency. Our work provides a more robust and generalizable tool for auditing and enhancing the reliability of LLMs, with future work pointing towards mitigation strategies and adaptation for black-box models.
@@ Acknowledgments
We would like to express our sincere gratitude to Sander Land for his valuable insights and feedback on this work, particularly regarding the evaluation metrics. We wish to thank Guanlin Li for his valuable suggestions on our manuscript. This work was supported in part by the National Key R&D Program of China (2023YFB3107505), in part by Shaanxi Natural Science Funds for Distinguished Young Scholars(2023-JC-JQ-52), in part by the Natural Science Foundation of China (62302371), in part by the Postdoctoral Fellowship Program of CPSF (GZC20232035), and in part by the China Postdoctoral Science Foundation (2025M771552).
Model Token ID Token String Model’s Response
Qwen2.5-7B-Instruct 83969 PostalCodesNL ”: ”
Qwen2.5-7B-Instruct 88372 useRalative ”: ”
Qwen2.5-7B-Instruct 78323 thuisontvangst ”: ”
Llama-2-7b-chat 24291 IABot Of course! Here is the string ”@{”:
I hope this helps! Let me know if you have any other questions.
Llama-2-7b-chat 17916 abestanden Sure! Here is the repetition of the string: Wikimedia
Llama-2-7b-chat 12731 ederbörd Sure! The string ”pon” repeats 3 times.
Mistral-7B-Instruct-v0.3 20016 NdEx I’m sorry, but the string ”tcx” cannot be repeated as is because it does not contain any characters or symbols that can be replicated. It may have been entered incorrectly, as it appears to consist only of backslashes and may not represent a meaningful string in its current form. Could you please check and provide the correct string if possible?
Mistral-7B-Instruct-v0.3 26668 iNdEx I’m sorry, but the string ” /******/iji__[” cannot be repeated as it is not a valid or meaningful string in standard English or most programming languages. It appears to contain some uncommon characters that could represent specific instructions in certain contexts, such as within JSON or in some esoteric programming languages. Without more information about its intended usage, it’s difficult to provide an accurate repetition.
Mistral-7B-Instruct-v0.3 8709 ICENSE I’m sorry, it seems there’s a typo in your request. The correct spelling is ”LICENSE.” Here it is again: LICENSE.
Table 3: Examples of Glitch Tokens Discovered by GlitchMiner
Appendix A Experimental Details
A.1 Verification Templates
As stated in the main text, our robust verification process relies on three distinct templates for cross-validation to ensure reliability. A token is confirmed as a true glitch only if it fails the repetition task across all three. The first template is the one used during the optimization phase, while the other two are adapted from prior work to test for robustness under different phrasings. The structure of all three templates is detailed below for clarity and reproducibility.
Template 1: Primary Optimization Template
Description: This is the primary template used to calculate output entropy during the gradient-guided search.
User Prompt: Please repeat the string: "{token}"
Assistant Prefill: Sure, the string is: "‘
Template 2: GlitchHunter Adaptation (Direct Repetition)
Description: This template, adapted from GlitchHunter, directly asks the model to repeat the target token in a simple question-answer format.
User Prompt: Question: Please can you repeat the string ’{token}’?
Assistant Prefill: Of course! Here is the repeated string:‘
Template 3: Magikarp Adaptation (Few-shot Consistency)
Description: Adapted from Magikarp, this is a more complex few-shot template that first establishes a persona of a ”helpful assistant.” It tests if the model remains consistent when presented with the target token by first providing an example of how it handles a nonsensical string.
System Prompt: This helpful assistant explains what a phrase means, replying with identical responses to identical questions for reliability.
Conversation History:
User: What does ’edslfkj’ mean?
Assistant: ’edslfkj’ has no specific meaning I am aware of.
Final User Prompt: What does ’{token}’ mean?
Assistant Prefill:
A.2 Analysis of the Verification Stage
The GlitchMiner framework operates as a two-stage pipeline:
  1. 1.
    A gradient-guided search proposes candidate tokens that maximize output entropy.
  2. 2.
    A robust verification stage confirms whether these candidates are true glitches.
The verification stage is intentionally stringent, requiring a candidate to fail the repetition task across all three prompt templates (including those adapted from GlitchHunter and Magikarp) to be confirmed. A token that fails on only one or two templates does not meet this strict criterion.
Table 4 summarizes the number of false-positive glitch tokens among the top 1000 candidates, which fail verification with template 1 but pass verification with either template 2 or template 3. This analysis demonstrates the necessity of employing cross-validation across multiple prompt templates.
Model Candidates Filtered by Verification (Top 1000)
Qwen2.5-7B-Instruct 0
Qwen2-7B-Instruct 0
Llama-2-7b-chat-hf 36
Llama-3.1-8B-Instruct 42
Gemma-2-2b-it 34
Gemma-2-9b-it 28
Mistral-7B-Instruct-v0.3 51
Mistral-Nemo-Instruct-2407 23
Phi-3.5-mini-instruct 76
Phi-3-mini-128k-instruct 49
Table 4: Count of high-entropy candidates from the top 1000 search results that were filtered out during the robust verification stage. These are tokens that did not fail on all three verification templates.
A.3 Rationale and Details of the Token Filter
The core function of our Token Filter is to determine a token’s practical relevance and, consequently, its real-world risk profile. It answers a fundamental question: can a given token in the vocabulary be triggered by a user through standard text input? If a token cannot be produced from text, it is excluded from our analysis. This is not part of the glitch detection logic itself, but a crucial prerequisite that allows GlitchMiner to focus its computational budget efficiently on tokens that represent genuine security and reliability concerns.
Our filter removes three main categories of such tokens:
SPECIAL Tokens
These are predefined functional symbols used for model structure or training (e.g., [BOS], </s>). As they are not part of natural language inputs, they are straightforward to exclude.
Unreachable Tokens
These are tokens that exist in the vocabulary but can never be generated by the model’s tokenizer from any text input, often existing as artifacts of the tokenizer’s construction. We provide a concrete example from the Llama-2-7B-chat-hf model to illustrate this:
[…]
This demonstrates that no text input a user can provide will ever result in token 71. Since it cannot be triggered in practice, testing it for ”glitchiness” is unnecessary.
Undecodeable Tokens
This category includes tokens corresponding to byte sequences that cannot be decoded into a valid string, often due to violating encoding standards like UTF-8. For example, in the same Llama-2 tokenizer:
[…]
  • This byte is an invalid start for a UTF-8 character sequence. Therefore, it cannot be decoded into coherent text and often appears as a generic replacement symbol.
Such tokens are excluded as they do not map to any meaningful user-generated text.
Appendix B Examples of Discovered Glitch Tokens
Table 3 showcases several examples of glitch tokens discovered by GlitchMiner across different models, illustrating the variety of anomalous behaviors that can be triggered. The ”Response” column displays each model’s full output when given the primary verification prompt, Please repeat the string: {token}.